OpenStudy (anonymous):

Solve. 3/x-1 +2/x=4 A) 0.25, 2 B)1/2, 2 C)1/2 , 1 D) .25 , 1

hero (hero):

\(\large\frac{3}{x-1} + \frac{2}{x} = 4 \)

OpenStudy (anonymous):



OpenStudy (anonymous):

yeah

hero (hero):

First write 4 as a fraction: \(\frac{4}{1}\)

OpenStudy (anonymous):

alright

hero (hero):

Next multiply \(\frac{4}{1}\) by \(\frac{x}{x}\)

hero (hero):

Afterwards, subtract \(\frac{2}{x}\) from both sides

hero (hero):

You should have: \(\frac{3}{x-1} = \frac{4x - 2}{x}\)

hero (hero):

Now cross multiply, and continue solving for x

hero (hero):

Have you cross multiplied yet?

OpenStudy (anonymous):

i dont understand how to cross multiply it

OpenStudy (anonymous):

i tried a couple different ways but it was confusing me and sorry it took so long had to help a family member with something

hero (hero):

You mean you don't know how to multiply (3)(x) = (4x-2)(x-1) ?

hero (hero):

I just cross multiplied it for you

hero (hero):

All you have to do is simplify it from there

OpenStudy (anonymous):

the (3)(x) is easy but the other half confuses me

hero (hero):

3x = x(4x-2) - 1(4x-2) 3x = 4x^2 - 2x - 4x + 2 3x = 4x^2 - 6x + 2 0 = 4x^2 - 6x - 3x + 2 0 = 4x^2 - 9x + 2 0 = 4x^2 -8x - 1x + 2 0 = 4x(x-2) -1(x-2) 0 = (x-2)(4x - 1) Continue solving for x

OpenStudy (anonymous):

i substitued the answers from the multiple choice answers and ended up with the answer , thanks

hero (hero):

lol, you never said you had any multiple choices.

hero (hero):

x = 2, 1/4